# A Polylogarithmic Competitive Algorithm for Stochastic Online Sorting and TSP

### v2

@@ Abstract
In Online Sorting, an array of $n$ initially empty cells is given. At each time step $t$, an element $x_{t}\in[0,1]$ arrives and must be placed irrevocably into an empty cell without any knowledge of future arrivals. We aim to minimize the sum of absolute differences between pairs of elements placed in consecutive array cells, seeking an online placement strategy that results in a final array close to a sorted one. An interesting multidimensional generalization, a.k.a. the Online Travelling Salesperson Problem, arises when the request sequence consists of points in the $d$-dimensional unit cube and the objective is to minimize the sum of euclidean distances between points in consecutive cells. Motivated by the recent work of (Abrahamsen, Bercea, Beretta, Klausen and Kozma; ESA 2024), we consider the stochastic version of Online Sorting (resp. Online TSP), where each element (resp. point) $x_{t}$ is an i.i.d. sample from the uniform distribution on $[0,1]$ (resp. $0,1]^{d}$). By carefully decomposing the request sequence into a hierarchy of balls-into-bins instances, where the balls to bins ratio is large enough so that bin occupancy is sharply concentrated around its mean and small enough so that we can efficiently deal with the elements placed in the same bin, we obtain an online algorithm that approximates the optimal cost within a factor of $O(\log^{2}n)$ with high probability. Our result comprises an exponential improvement on the previously best known competitive ratio of $\tilde{O}(n^{1/4})$ for Stochastic Online Sorting due to (Abrahamsen et al.; ESA 2024) and $O(\sqrt{n})$ for (adversarial) Online TSP due to (Bertram, ESA 2025).
@@ 1 Introduction
In the Online Sorting Problem we are given a sequence of $n$ real numbers $x_{1},x_{2},\ldots,x_{n}\in[0,1]$, revealed one by one in an online fashion. An array $A$ of length $n$ is initially empty. Denote by $A[1],A[2],\ldots,A[n]$ its cells. Upon the arrival of each element $x_{j}$, the algorithm must immediately and irrevocably assign it to an empty cell of $A$. After all $n$ elements have been placed, the cost is defined as the total variation between consecutive entries i.e., $\text{Cost}(A)=\sum_{i=1}^{n-1}|{|A[i+1]-A[i]}|$. The objective is to minimize this cost. The problem was introduced by Aamand, Abrahamsen, Beretta, and Kleist [AABK23] as a technical tool for proving lower bounds in online strip packing, bin packing, and perimeter packing. They studied the adversarial setting, where the sequence is chosen by an adaptive adversary, and designed an $O(\sqrt{n})$-competitive algorithm along with a matching lower bound for deterministic algorithms. Later, Abrahamsen, Bercea, Beretta, Klausen, and Kozma [ABB+24] showed that even randomized algorithms cannot asymptotically improve this guarantee in the worst case.
@@
Abrahamsen et al. [ABB+24] further generalized the problem by increasing its dimensionality, introducing the Online Travelling Salesperson Problem. Here the input elements $x_{1},\dots,x_{n}$ are points in $0,1]^{d}$, and the objective is to minimize the total Euclidean distance between consecutive points in the array. In the adversarial setting, they gave a dimension-dependent algorithm with competitive ratio $\sqrt{d}\cdot 2^{d}\cdot O(\sqrt{n\log n})$. Shortly after, Bertram [Ber25] showed that dimension-independent bounds are achievable in the adversarial model presenting an $O(\sqrt{n})$-competitive algorithm. To the best of our knowledge, the competitive ratio of the stochastic version of this problem has not been studied.
@@ 1.2 Technical Overview
The first main idea that Abrahamsen et al. [ABB+24] introduced is to divide the array into buckets / bins, exploit the randomness of the input in order to assign each element to a bucket, and solve the problem separately in each bucket. Suppose that each bucket has $C$ elements, thus we have $n/C$ buckets and suppose bucket $i$ contains elements in $(\frac{i-1}{C},\frac{i}{C}]$. Then, assuming that each bucket will receive exactly the expected number of elements, i.e. $C$ elements, the cost of the algorithm is approximately $\textsc{Alg}=\text{number of buckets}\times\text{cost inside bucket}=O(\sqrt{C})$, assuming that the $O(\sqrt{n})$ adversarial algorithm of Aamand et al. [AABK23] was employed in each bucket.
Hence, smaller buckets reduce cost, but increase the chance of imbalance, while larger buckets stabilize the load at the expense of higher internal cost. This trade-off is central to their analysis. To handle the imbalances, a part of the array in the end is reserved, a.k.a. backyard, and its size depends on how large these imbalances are. Since elements inserted in the backyard are arbitrary the algorithm incurs cost equal to the square root of its size. Thus, the previous trade-off translates to a trade-off between the size of the backyard and $C$: the larger $C$ is, the smaller the backyard is and vice versa. Their first algorithm balances this trade-off and achieves $\widetilde{O}(n^{1/3})$ complexity. Recursively using their algorithm inside the buckets they improve their competitive ratio to $\widetilde{O}(n^{1/4})$.
Our approach to Stochastic Online Sorting can be naturally described through the lens of the balls-into-bins paradigm. More specifically, the algorithm proceeds by maintaining a partitioning of a certain part of $A$ into equally sized bins, while the elements arriving can be regarded as balls placed into bins. Each bin is associated with a certain subinterval of $[0,1]$, which determines if a new ball / element is placed into the particular bin. The subintervals associated with the bins form a partitioning of $[0,1]$ and are of equal length. Hence, since the elements are i.i.d samples from the uniform distribution on $[0,1]$, each new ball is placed uniformly at random (and independently of the other balls) in each bin.
@@
Our algorithm (see also Section 2 and Algorithm 1 for the details) maintains that all bins in the same phase are of equal size (and have equal probability of receiving a new element) and all bins created by the algorithm have size $\Theta(\log^{2}n)$ (though the bucket size may slightly vary between phases). In the first phase, we consider the first $n/2$ cells of $A$, which we partition into $n/(2\log^{2}n)$ buckets (we also partition $[0,1]$ into the same number of subintervals with equal length). The current phase ends when the first of its buckets becomes full. A standard concentration bound (Lemma 2.3) shows that w.h.p. bfthe first phase ends: (i) at least $(1-o(1))\frac{n}{2}$ elements are successfully placed into its bins; and (ii) every bin has received at least $(1-o(1))\log^{2}n$ elements. Using the (adversarial) $O(\sqrt{n})$-competitive algorithm of Aamand et al. [AABK23] to deal with the exact placement of the elements in the same bin we get that w.h.p. the total cost of the algorithm during its first phase is $O(\log n)$.
@@ 1.3 Other Related Work
Another interesting variant of the Online Sorting Problem is one where the size of the array $m$ is longer than the input sequence, i.e., $m>n$. This version was introduced by Aamand et al. [AABK23], who designed a deterministic $2^{\sqrt{\log n}\sqrt{\log\log n+\log(1/\epsilon)}}$-competitive algorithm when $m=(1+\epsilon)n$. They complemented this with a lower bound, showing that every deterministic algorithm with $m=\gamma n$ is at least $1/\gamma\cdot\Omega(\log n/\log\log n)$-competitive. Later, Azar et al. [APV25] and Nirjhor et al. [NW25], in independent and concurrent work, improved the upper bound, nearly resolving this variant.
@@
The connection between stochastic online sorting and hashing was already observed by Abrahamsen et al. [ABB+24]. Two particularly relevant examples are Filter Hashing, introduced by Fotakis et al. [FPSS03], and Transactional Multi-Writer Cuckoo Hashing, proposed by Kuszmaul [Kus16]. Both hashing schemes employ a multi-layered structure where each subsequent layer is tasked with handling the imbalances of the previous layers.
@@ 2 General Algorithmic Framework
Before we present our algorithm, we introduce some notations. For an array $A$ we denote by $A[s:t]$ the subarray of $A$ starting at $s$ and ending at $t$. We also adopt a slight abuse of notation by using $A$ to refer both to the array itself and to its length, with the intended meaning being clear from the context. We also define by $[k]$ the set $\{1,2,\ldots,k\}$. We call the subarray of a subarray of array $A$, a sub-subarray (or bucket).
@@ 2.1 Algorithm Description
  • $\texttt{DomainPartitioning}\left(\mathcal{D},\mathcal{S},\ell\right)$: This procedure takes as input a distribution $\mathcal{D}$ over a domain $\mathcal{S}$ and a positive integer $\ell$. It outputs subsets $\bigl\{T^{(i)}_{j}\subseteq\mathcal{S}:i\in[\ell],\;j\in[2^{i}]\bigr\},$ collectively denoted by $\mathcal{T}$. The collection $\mathcal{T}$ satisfies, for every level $i\in[\ell]$:
@@
      Covering: $\bigcup_{j\in[2^{i}]}T^{(i)}_{j}=\mathcal{S}.$
@@
      Disjointness: The sets $\{T^{(i)}_{j}\}_{j\in[2^{i}]}$ are pairwise disjoint.
@@
      Equal Mass: $\mathbb{P}_{x\sim\mathcal{D}}\!\left[x\in T^{(i)}_{j}\right]\text{ is the same for all }j\in[2^{i}].$
@@
  • $\texttt{index}(\mathcal{T},i)$: this function takes as input a set $\mathcal{T}$ outputted by DomainPartitioning and a positive integer $i$ and outputs the unique index $k$ such that $x\in T^{(i)}_{k}$.
  • empty($a$): takes as input a (sub-)subarray $a$ of $A$ and returns the number of empty cells in $a$ at the current point of execution.
@@
For a subarray $i$, we define as $N_{i}\coloneqq\texttt{empty}(A_{i})$ at the time of the overflow of $A_{i}$.
@@ 1 $i\leftarrow 0;$
3 $\mathcal{T}\leftarrow\texttt{DomainPartitioning}\left(\mathcal{D},\mathcal{S},K\right);$
@@ 13    $x\leftarrow\texttt{receive\_sample}(\mathcal{D})$;
14    $k_{1}\leftarrow\texttt{index}(\mathcal{T},i-1)$, $k_{2}\leftarrow\texttt{index}(\mathcal{T},i)$;
@@ 18       $\texttt{InBucketPlacement}\left(x,a^{(i)}_{k_{2}}\right)$;
In order to handle overflows from subarray $A_{1}$ we allocate a subarray $A_{2}$ of size $\left\lfloor\frac{n}{4}\right\rfloor$ next to $A_{1}$. Subarray $A_{2}$ is divided into $K_{2}=K_{1}/2=2^{\ell-1}$ buckets. Regarding bucket sizes, crucially, the size of each bucket of $A_{2}$ is not the same, as discussed in technical overview. To gain some intuition for the bucket sizes set in Algorithm˜1, we consider the following example.
@@
Suppose for simplicity’s sake, that we are on the classic online sorting case and subarray $A_{1}$ has $100$ cells and is divided into $K=K_{1}=4$ buckets, with capacity $C_{1}=25$ each, such that $T^{(1)}_{1}=(0,0.25),\ T^{(1)}_{2}=(0.25,0.5),\ T^{(1)}_{3}=(0.5,0.75),\ T^{(1)}_{4}=(0.75,1)$ respectively. Assume that bucket $a^{(1)}_{1}$ is overflowed and the number of elements in each bucket are $(25,20,10,15)$ at the time exactly before the overflow. Then, the up to this point empty, subarray $A_{2}$ will be brought into action to take care of the overflows of subarray $A_{1}$. By design, $A_{2}$ has $50$ cells and $K_{2}=K_{1}/2=2$ buckets that handle elements $T^{(2)}_{1}=(0,0.5),\ T^{(2)}_{2}=(0.5,1)$ respectively. Our goal is to set the sizes of each bucket in $A_{2}$ in a way that ensure that an overflow is occurred relatively late, for instance after $A_{1}$ is completely filled. If we allocate $25$ cells to each bucket in $A_{2}$, then we expect the first one to overflow significantly faster than the second one since in $A_{1}$ there are less empty cells for elements lying in $(0,0.5)$ than for elements lying in $(0.5,1)$. Thus, in order to handle the imbalances in $A_{1}$ we will allocate more cells in the first bucket in $A_{2}$. The total number of empty cells in $A_{1}$ is $\texttt{empty}(A_{1})=0+5+15+10=30$ and since we add $50$ cells with $A_{2}$ we have in total $80$ cells—this is the size of the second balls-into-bins instance. Specifically, we can consider that the instance has $2$ bins that handle elements from $(0,0.5)$ and $(0.5,1)$ respectively. The first bin contains the first two buckets of $A_{1}$ and the first bucket of $A_{2}$. The second bin contains the last two buckets of $A_{1}$ and the second bucket of $A_{2}$. We would like the capacities of the bins to be close enough in order to have a balanced balls-into-bins instance, thus equal to $\frac{\texttt{empty}(A_{1})+A_{2}}{K_{2}}=40$. Thus, we allocate $c^{(2)}_{1}=40-\texttt{empty}(a^{(1)}_{1})-\texttt{empty}(a^{(1)}_{2})=35$ cells for the first bucket in $A_{2}$ and $c^{(2)}_{2}=40-\texttt{empty}(a^{(1)}_{3})-\texttt{empty}(a^{(1)}_{4})=15$ cells for the second one. As a result, we obtain a new balanced balls-into-bins instance: the bins have equal capacity, so the expected overflow occurs later than it would if $A_{2}$ were simply divided into two equal-sized buckets.
@@
$C_{2}=\frac{\text{size of instance}}{\text{\# of bins}}=\frac{A_{2}+\texttt{empty}(A_{1})}{K_{2}}=\Theta(\log^{2}n)$
@@
$\sum_{i=1}^{R-1}\frac{n}{2^{i}}\geq n-100\log^{2}n\Rightarrow n(1-2^{-R})\geq n-100\log^{2}n\Rightarrow R\geq\log\left(\frac{n}{100\log^{2}n}\right)$ (2)
and since $n-\sum_{i=1}^{R-2}\left\lfloor\frac{n}{2^{i}}\right\rfloor>100\log^{2}n$ we get that $R\leq 1+\log\left(\frac{n}{100\log^{2}n}\right)$, thus $R=\Theta(\log\frac{n}{\log^{2}n})$. Also, the size of $A_{R-1}$ is $n/2^{R-1}=\Theta(\log^{2}n)$ and for $B$ it holds that $B\geq n-n(1-2^{-R})=\Theta(\log^{2}n)$, thus $B$ and $A_{T-1}$ are asymptotically tight, preserving the invariant.

### v3

@@ Abstract
In Online Sorting, an array of $n$ initially empty cells is given. At each time step $t$, an element $x_{t}\in[0,1]$ arrives and must be irrevocably placed in an empty cell without knowledge of future arrivals. We aim to minimize the sum of absolute differences between pairs of elements placed in consecutive array cells, seeking an online placement strategy that results in a final array close to a sorted one. An interesting multidimensional generalization, referred to as the Online Traveling Salesperson Problem, arises when the request sequence consists of points in the $d$-dimensional unit cube and the objective is to minimize the sum of Euclidean distances between points in consecutive cells. Motivated by the recent work of (Abrahamsen, Bercea, Beretta, Klausen and Kozma; ESA 2024), we consider the stochastic version of Online Sorting (resp. Online TSP), where each element (resp. point) $x_{t}$ is an i.i.d. sample from the uniform distribution on $[0,1]$ (resp. $0,1]^{d}$). By carefully decomposing the request sequence into a hierarchy of balls-into-bins instances, where the balls to bins ratio is large enough so that bin occupancy is sharply concentrated around its mean and small enough so that we can efficiently deal with the elements placed in the same bin, we obtain an online algorithm that approximates the optimal cost within a factor of $O(\log^{2}n)$ with high probability. Our result comprises an exponential improvement over the previously best known competitive ratio of $\tilde{O}(n^{1/4})$ for Stochastic Online Sorting due to (Abrahamsen et al.; ESA 2024) and $O(\sqrt{n})$ for (adversarial) Online TSP due to (Bertram, ESA 2025).
@@ 1 Introduction
In the Online Sorting Problem we are given a sequence of $n$ real numbers $x_{1},x_{2},\ldots,x_{n}\in[0,1]$, revealed one by one in an online fashion. An array $A$ of length $n$ is initially empty. Denote by $A[1],A[2],\ldots,A[n]$ its cells. Upon the arrival of each element $x_{j}$, the algorithm must immediately and irrevocably assign it to an empty cell of $A$. After all $n$ elements have been placed, the cost is defined as the total variation between consecutive entries i.e., $\text{Cost}(A)=\sum_{i=1}^{n-1}|{A[i+1]-A[i]}|$. The objective is to minimize this cost. The problem was introduced by Aamand, Abrahamsen, Beretta, and Kleist [AAB+23] as a technical tool for proving lower bounds in online strip packing, bin packing, and perimeter packing. They studied the adversarial setting, where the sequence is chosen by an adaptive adversary, and designed an $O(\sqrt{n})$-competitive algorithm along with a matching lower bound for deterministic algorithms. Later, Abrahamsen, Bercea, Beretta, Klausen, and Kozma [ABB+24] showed that even randomized algorithms cannot asymptotically improve this guarantee in the worst case.
@@
Abrahamsen et al. [ABB+24] further generalized the problem by increasing its dimensionality, introducing the Online Traveling Salesperson Problem. Here the input elements $x_{1},\dots,x_{n}$ are points in $0,1]^{d}$, and the objective is to minimize the total Euclidean distance between consecutive points in the array. In the adversarial setting, they gave a dimension-dependent algorithm with competitive ratio $\sqrt{d}\cdot 2^{d}\cdot O(\sqrt{n\log n})$. Shortly after, Bertram [BER25] showed that dimension-independent bounds are achievable in the adversarial model presenting an $O(\sqrt{n})$-competitive algorithm. To the best of our knowledge, the competitive ratio of the stochastic version of this problem has not been studied.
@@ 1.2 Technical Overview
The first main idea that Abrahamsen et al. [ABB+24] introduced is to divide the array into buckets / bins, exploit the randomness of the input in order to assign each element to a bucket, and solve the problem separately in each bucket. Suppose that each bucket has $C$ elements; thus, we have $n/C$ buckets, and the bucket $i$ contains elements in $((i-1)\frac{C}{n},i\frac{C}{n}]$. Then, assuming that each bucket will receive exactly the expected number of elements, i.e. $C$ elements, the cost of the algorithm is approximately $\textsc{Alg}=\text{number of buckets}\times\text{cost inside bucket}=O(\sqrt{C})$, assuming that the $O(\sqrt{n})$ adversarial algorithm of Aamand et al. [AAB+23] was employed in each bucket.
Hence, smaller buckets reduce cost but increase the chance of imbalance, while larger buckets stabilize the load at the expense of higher internal cost. This trade-off is central to their analysis. To handle the imbalances, a part of the array in the end is reserved, also called backyard, and its size depends on how large these imbalances are. Since elements inserted in the backyard are arbitrary, the algorithm incurs a cost equal to the square root of its size. Thus, the previous trade-off translates into a trade-off between the size of the backyard and $C$: the larger $C$ is, the smaller the backyard and vice versa. Their first algorithm balances this trade-off and achieves $\widetilde{O}(n^{1/3})$ complexity. Recursively using their algorithm inside the buckets, they improve their competitive ratio to $\widetilde{O}(n^{1/4})$.
Our approach to Stochastic Online Sorting can be naturally described through the lens of the balls-into-bins paradigm. More specifically, the algorithm proceeds by maintaining a partitioning of a certain part of $A$ into equally sized bins, while the arriving elements can be regarded as balls placed into bins. Each bin is associated with a certain subinterval of $[0,1]$, which determines if a new ball / element is placed into the particular bin. The subintervals associated with the bins form a partitioning of $[0,1]$ and are of equal length. Hence, since the elements are i.i.d. samples from the uniform distribution on $[0,1]$, each new ball is placed uniformly at random (and independently of the other balls) in each bin.
@@
Our algorithm (see also Section 2 and Algorithm 1 for the details) maintains that all bins in the same phase are of equal size (and have equal probability of receiving a new element) and all bins created by the algorithm have size $\Theta(\log^{2}n)$ (though the bucket size may slightly vary between phases). In the first phase, we consider the first $n/2$ cells of $A$, which we partition into $n/(2\log^{2}n)$ buckets (we also partition $[0,1]$ into the same number of subintervals with equal length). The current phase ends when the first of its buckets becomes full. A standard concentration bound (Lemma 2.3) shows that w.h.p. before the first phase ends: (i) at least $(1-o(1))\frac{n}{2}$ elements are successfully placed into its bins; and (ii) every bin has received at least $(1-o(1))\log^{2}n$ elements. Using the (adversarial) $O(\sqrt{n})$-competitive algorithm of Aamand et al. [AAB+23] to deal with the exact placement of the elements in the same bin we get that w.h.p. the total cost of the algorithm during its first phase is $O(\log n)$.
@@ 1.3 Other Related Work
Another interesting variant of the Online Sorting Problem is one where the size of the array $m$ is longer than the input sequence, i.e., $m>n$. This version was introduced by Aamand et al. [AAB+23], who designed a deterministic $2^{\sqrt{\log n}\sqrt{\log\log n+\log(1/\epsilon)}}$-competitive algorithm when $m=(1+\epsilon)n$. They complemented this with a lower bound, showing that every deterministic algorithm with $m=\gamma n$ is at least $1/\gamma\cdot\Omega(\log n/\log\log n)$-competitive. Later, Azar et al. [APV25] and, independently and concurrently, Nirjhor et al. [NW25] improved the upper bound, with the former nearly resolving this variant.
@@
The connection between stochastic online sorting and hashing was already observed by Abrahamsen et al. [ABB+24]. Two particularly relevant examples are Filter Hashing, introduced by Fotakis et al. [FPS+03], and Transactional Multi-Writer Cuckoo Hashing, proposed by Kuszmaul [KUS16]. Both hashing schemes employ a multi-layered structure where each subsequent layer is tasked with handling the imbalances of the previous layers.
@@ 2 General Algorithmic Framework
Before we present our algorithm, we introduce some notation. For an array $A$ we denote by $A[s:t]$ the subarray of $A$ starting at $s$ and ending at $t$. We also adopt a slight abuse of notation by using $A$ to refer both to the array itself and to its length, with the intended meaning being clear from the context. We also define by $[k]$ the set $\{1,2,\ldots,k\}$. We call a subarray of a subarray of $A$ a sub-subarray (or bucket).
@@ 2.1 Algorithm Description
  • $\texttt{DomainPartitioning}\left(\mathcal{D},\mathcal{S},\ell\right)$: This procedure takes as input a distribution $\mathcal{D}$ over a domain $\mathcal{S}$ and a positive integer $\ell$. It outputs subsets $\bigl\{T^{(i)}_{j}\subseteq\mathcal{S}:i\in[\ell+1],\;j\in[2^{\ell-i+1}]\bigr\},$ collectively denoted by $\mathcal{T}$. The collection $\mathcal{T}$ satisfies, for every level $i\in[\ell+1]$:
@@
      Covering: $\bigcup_{j\in[2^{\ell-i+1}]}T^{(i)}_{j}=\mathcal{S}.$
@@
      Disjointness: The sets $\{T^{(i)}_{j}\}_{j\in[2^{\ell-i+1}]}$ are pairwise disjoint.
@@
      Equal Mass: $\mathbb{P}_{x\sim\mathcal{D}}\!\left[x\in T^{(i)}_{j}\right]\text{ is the same for all }j\in[2^{\ell-i+1}].$
@@
  • $\texttt{index}(\mathcal{T},i,x)$: This function takes as input a set $\mathcal{T}$ outputted by DomainPartitioning, a positive integer $i$ and an input element $x$ and outputs the unique index $k$ such that $x\in T^{(i)}_{k}$.
  • empty($a$): This function takes as input a (sub-)subarray $a$ of $A$ and returns the number of empty cells in $a$ at the current point of execution.
@@
For a subarray $A_{i}$, let $N_{i}\coloneqq\texttt{empty}(A_{i})$ at the moment when the algorithm transitions from phase $i$ to phase $i+1$, i.e., when some bucket of $A_{i}$ becomes full and triggers an overflow.
@@ 1 $i\leftarrow 0;$
3 $\mathcal{T}\leftarrow\texttt{DomainPartitioning}\left(\mathcal{D},\mathcal{S},\ell\right);$
@@ 13    $x\leftarrow\texttt{receive\_sample}(\mathcal{D})$;
14    $k_{1}\leftarrow\texttt{index}(\mathcal{T},i-1,x)$, $k_{2}\leftarrow\texttt{index}(\mathcal{T},i,x)$;
@@ 18       $\texttt{InBucketPlacement}\left(x,a^{(i)}_{k_{2}}\right)$;
In order to handle overflows from subarray $A_{1}$ we allocate a subarray $A_{2}$ of size $\left\lfloor\frac{n}{4}\right\rfloor$ next to $A_{1}$. Subarray $A_{2}$ is divided into $K_{2}=K_{1}/2=2^{\ell-1}$ buckets. Regarding bucket sizes, crucially, the size of each bucket of $A_{2}$ is not the same, as discussed in the technical overview. To gain some intuition for the bucket sizes set in Algorithm˜1, we consider the following example.
@@
Suppose for simplicity’s sake, that we are on the classic online sorting case and subarray $A_{1}$ has $100$ cells and is divided into $K=K_{1}=4$ buckets, with capacity $C_{1}=25$ each, such that $T^{(1)}_{1}=(0,0.25),\ T^{(1)}_{2}=(0.25,0.5),\ T^{(1)}_{3}=(0.5,0.75),\ T^{(1)}_{4}=(0.75,1)$ respectively. Assume that bucket $a^{(1)}_{1}$ is overflowed and the number of elements in each bucket are $(25,20,10,15)$ at the time exactly before the overflow. Then, the up to this point empty, subarray $A_{2}$ will be brought into action to take care of the overflows of subarray $A_{1}$. By design, $A_{2}$ has $50$ cells and $K_{2}=K_{1}/2=2$ buckets that handle elements $T^{(2)}_{1}=(0,0.5),\ T^{(2)}_{2}=(0.5,1)$ respectively. If we allocate $25$ cells to each bucket in $A_{2}$, then we expect the first one to overflow significantly faster than the second one since in $A_{1}$ there are less empty cells for elements lying in $(0,0.5)$ than for elements lying in $(0.5,1)$. Thus, in order to handle the imbalances in $A_{1}$ we will allocate more cells in the first bucket in $A_{2}$. The total number of empty cells in $A_{1}$ is $\texttt{empty}(A_{1})=0+5+15+10=30$ and since we add $50$ cells with $A_{2}$ we have in total $80$ cells—this is the size of the second balls-into-bins instance. Specifically, we can consider that the instance has $2$ bins that handle elements from $(0,0.5)$ and $(0.5,1)$ respectively. The first bin contains the first two buckets of $A_{1}$ and the first bucket of $A_{2}$. The second bin contains the last two buckets of $A_{1}$ and the second bucket of $A_{2}$. We would like the capacities of the bins to be close enough in order to have a balanced balls-into-bins instance, thus equal to $\frac{\texttt{empty}(A_{1})+A_{2}}{K_{2}}=40$. Thus, we allocate $c^{(2)}_{1}=40-\texttt{empty}(a^{(1)}_{1})-\texttt{empty}(a^{(1)}_{2})=35$ cells for the first bucket in $A_{2}$ and $c^{(2)}_{2}=40-\texttt{empty}(a^{(1)}_{3})-\texttt{empty}(a^{(1)}_{4})=15$ cells for the second one. As a result, we obtain a new balanced balls-into-bins instance: the bins have equal capacity, so the expected overflow occurs later than it would if $A_{2}$ were simply divided into two equal-sized buckets.
@@
$C_{2}=\frac{\text{size of instance}}{\text{number of bins}}=\frac{A_{2}+\texttt{empty}(A_{1})}{K_{2}}=\Theta(\log^{2}n)$
@@
$\sum_{i=1}^{R-1}\frac{n}{2^{i}}\geq n-100\log^{2}n\Rightarrow n(1-2^{-R})\geq n-100\log^{2}n\Rightarrow R\geq\log\left(\frac{n}{100\log^{2}n}\right)$
and since $n-\sum_{i=1}^{R-2}\left\lfloor\frac{n}{2^{i}}\right\rfloor>100\log^{2}n$ we get that $R\leq 1+\log\left(\frac{n}{100\log^{2}n}\right)$, thus $R=\Theta(\log\frac{n}{\log^{2}n})$. Also, the size of $A_{R-1}$ is $n/2^{R-1}=\Theta(\log^{2}n)$ and for $B$ it holds that $B\geq n-n(1-2^{-R})=\Theta(\log^{2}n)$, thus $B$ and $A_{R-1}$ are asymptotically tight, preserving the invariant.